The sum of the digits of a two-digit number is 13. the value of the number is 2 more than 7 times the units digit

Answer by lwsshak3(11628)

The sum of the digits of a two-digit number is 13. the value of the number is 2 more than 7 times the units digit
***
let u=units digit
let t=tens digit
u+t=13
t=13-u
..
10t+u=7u+2
10(13-u)+u=7u+2
130-10u+u=7u+2
16u=128
u=8
t=5
number=58